A person's blood has a pH of 7.408 and a bicarbonate concentration of 26.4. Approximate the concentration of carbonic acid in the person's blood. We were given this formula: pH= 6.1 + log10(x/y) where x is the concentration of bicarbonate and y is the concentration of carbonic acid.
Answer by Nate(3500)

First, we have to shape the equation for y:
pH= 6.1 + log10(x/y)
pH - 6.1 = log10(x/y)
(pH - 6.1)/log10 = x/y
((pH - 6.1)/log10)y = x
y = (x)/((pH - 6.1)/log10)
Now, lets plug in the numbers:
y = (x)/((pH - 6.1)/log10)
y = (26.4)/((7.408 - 6.1)/(log10))
y = (26.4) / (1.308)
y = (26400) / (1308)